kinetic energy of a particle moving in a straight line varies with time as k =4t2 . the force acting on the particle

Grade:11

1 Answers

Mallikarjun Maram
8 years ago
Given that K = 4 t2, then linear momentum P = \sqrt {2mK} = \sqrt {2 m 4 t ^ 2} = \sqrt {8m} \ t.  Hence force is given by F = \frac {dP}{dt} = \sqrt {8m} N, where m = mass of the particle.